OpenStudy (anonymous):

I really need the answer! Find parametric equations for the curve y = x + 1, x ≥ 0, by using as parameter the distance of the point (x,y) from the origin.

OpenStudy (anonymous):

I believe you'd start off by letting the new parameter be \(d\) and set it equal to \[d=\sqrt{x^2+y^2}\] Then, substitute \(y=x+1\): \[d=\sqrt{x^2+(x+1)^2}\\ d=\sqrt{x^2+x^2+2x+1}\\ d=\sqrt{2x^2+2x+1}\] Complete the square, you get \[d=\sqrt{2\left(x^2+x+\frac{1}{2}\right)}\\ d=\sqrt{2\left(x^2+x+\frac{1}{4}+\frac{1}{4}\right)}\\ d=\sqrt{2\left(\left(x+\frac{1}{2}\right)^2+\frac{1}{4}\right)}\\ d=\sqrt{2\left(x+\frac{1}{2}\right)^2+\frac{1}{2}}\] Now solve for x: \[d^2=2\left(x+\frac{1}{2}\right)^2+\frac{1}{2}\\ d^2-\frac{1}{2}=2\left(x+\frac{1}{2}\right)^2\\ \frac{d^2}{2}-\frac{1}{4}=\left(x+\frac{1}{2}\right)^2\\ \pm \sqrt{\frac{d^2}{2}-\frac{1}{4}}=x+\frac{1}{2}\\ \pm \sqrt{\frac{d^2}{2}-\frac{1}{4}}-\frac{1}{2}=x\] I think you can drop the plus or minus and only use the positive root, due to its symmetry, but I'm not 100% sure. Also, since x is negative for \(d<1\), you'll have \(d\ge1\) as a restriction. Here's a plot of the equation \(x=x(d)\): http://www.wolframalpha.com/input/?i=Plot%5B%7BSqrt%5Bd%5E2%2F2+-+1%2F4%5D-1%2F2%2C+-Sqrt%5Bd%5E2%2F2+-+1%2F4%5D-1%2F2%7D%2C%7Bd%2C-2%2C2%7D%5D

OpenStudy (anonymous):

Actually, forget what I said about the negative root. You have to drop it because you have negative values of x. Only use the positive root. Also, that thing I said about symmetry doesn't apply. I'm not sure what I was thinking.

OpenStudy (experimentx):

make use of parametric plot